### 26 Jan 2024 v1

@@ Abstract
Context:Stellar bars are non-axis symmetrical structures found in over $30$ per cent of massive disc galaxies in the local Universe. The environment could play a significant role in the probability of a spiral galaxy to develop or not a bar.
Aims:We investigate the influence of the environment on the evolution of barred and unbarred disc galaxies with a mass larger than $10^{10}\rm M_{\sun}$ from $z=1$ down to $z=0$, employing the TNG50 magnetic-hydrodynamical simulation.
Methods:We determine the fraction of barred galaxies that conserve their bar and the ones that lose their bar by $z=0$. We also estimate the fraction of unbarred galaxies at z=1 that develop a bar at later times. We study the merger histories and the distance of close companions for each category to understand the role of the environment in the evolution of these galaxies.
Results:We find that 49 per cent of $z=1$ disc galaxies undergoes a morphological transformation, transitioning into either a lenticular or spheroidal, while the other $51$ per cent retains the large disc shape. The morphological alteration is mostly influenced by the environment. Lenticular and spheroidal galaxies tend to exist in denser environments and have more frequent mergers compared to disc galaxies. We find that bars are stable after they have formed, as over half of the barred galaxies (60.2 per cent) retain the bar structure and have experienced fewer mergers compared to those galaxies that lose their bars (5.6 per cent). These latter ones start with weaker and shorter bars at $z=1$ and influenced by tidal interactions and are frequently observed in more populated areas. Additionally, our study reveals that less than 20 per cent of unbarred galaxies will never develop a bar and exhibit the quietest merger history. Unbarred galaxies that experience bar formation after $z=1$ exhibit more frequent instances of merging events. Furthermore, tidal interactions with a close companion may account for bar formation in at least one-third of the instances.
@@ 1 Introduction
Stellar bars, non-axis symmetrical structures that are found in the central regions of a majority of disc galaxies. They are present in over $30$ per cent of massive disc galaxies with a stellar mass of $10^{10}M_{\odot}$ in the local Universe (e.g., Sellwood & Wilkinson 1993; Masters et al. 2011; Gavazzi et al. 2015). These bars are believed to be drivers of the secular evolution of galaxies, as they are efficient in redistributing gas, stars, and dark matter towards the central regions of the galaxy (e.g., Debattista et al. 2004; Athanassoula et al. 2005).
The presence or absence of bars in galaxies can reveal crucial information about the assembly history of these galaxies. One effective way to investigate this is by studying the evolution of bar fractions as a function of various galaxy properties, such as stellar mass, gas content, or disc structure. Local Universe studies have shown that bars are more frequently found in massive, gas-poor, and red galaxies than in their blue, gas-rich, and star-forming counterparts (e.g., Barazza et al. 2008; Masters et al. 2012; Gavazzi et al. 2015; Consolandi 2016; Cervantes-Sodi 2017). Moreover, the redshift dependence of the bar fraction can provide additional insights into the evolution of bars and their host galaxies.
@@
The fact that the bar fraction is increasing or decreasing also encapsulates the lifetime of bar structures providing useful information about whether these structures are stables in time or just disappear through secular processes (e.g. Aguerri, Méndez-Abreu, & Corsini 2009) or through external processes (e.g. Zana et al. 2018b; Peschken & Lokas 2019; Méndez-Abreu, Costantin, & Kruk 2023). In this regard, there has been a lot of debate. The secular process of forming and dissolving a stellar bar can be complex and act over long timescales. Bars can be formed spontaneously via secular evolution as a result of global, non-axisymmetric instabilities (e.g. Athanassoula 2003; Mendez-Abreu et al. 2012; Gavazzi et al. 2015). On the other hand the combined effects of central mass concentrations (CMCs) and gravity torques could make galactic bars transient features with a life-time of 1-2 Gyr in typical Sb-Sc galaxies (Bournaud, Combes, & Semelin 2005).
@@
The formation of new stars can lead to feedback processes that release energy and momentum that can disrupt the bar structure and gradually weaken and dissolve it (e.g. Zana et al. 2019). Vertical resonances also can arise within the bar, causing stars to oscillate vertically within the bar. This can lead to the formation of a central mass concentration, which can destabilise the bar and cause it to dissolve (e.g. Martinez-Valpuesta, Shlosman, & Heller 2006).
@@
In the case of external processes, galactic mergers/interactions can play a significant role in the formation and disruption of stellar bars in galaxies (e.g. Elmegreen, Elmegreen, & Bellin 1990; Berentzen et al. 2004; Łokas et al. 2016; Zana et al. 2018a, b; Łokas 2019). The specific conditions for a galactic merger/interaction to form or dissolve a bar include the mass ratio of the merging galaxies, the phase of the bar and the orbit since the strength of the perturbation can be enough to affect the central part of the galaxy to form a bar (Elmegreen, Elmegreen, & Bellin 1990). The orbit of the merging galaxies can be another condition, a direct collision between the two galaxies can cause significant perturbations to form or dissolve bar (Berentzen et al. 2004). Moreover, prograde encounters (when the angular momenta of the two galaxies are aligned) have a much more dramatic effect on the galactic structure than retrograde ones (when the angular momenta point in opposite directions). Prograde encounters lead to the formation of long and narrow tidal arms and tidally induced bars in galaxies orbiting a bigger host, but not if they are on retrograde ones (Łokas 2018). The gas content of the merging galaxies also contributes to the dissolution of a bar. If the merging galaxies have a significant amount of gas, for instance, tidal interactions due to minor mergers can weaken the central stellar bar in the merger remnant (Ghosh et al. 2021).
@@
Tidal forces due to dense environments or close companions have also a deep impact on the formation and dissolution of bars. Łokas et al. (2014) and Łokas et al. (2016) study the formation of a bar in galaxy encounters with a Milky-way galaxy and dwarf galaxy embedded in dense environments using N-body simulations. The authors find that the characteristics of the bars undergo temporal variations and are subject to the influence of the magnitude of the tidal force encountered. The formation of bars in galaxies is observed to occur at earlier stages and exhibits greater strength and length in galaxies that experience a higher degree of tidal force exerted by the surrounding cluster.
@@
Until very recently witn zoom-in simulations, the external and internal processes could be studied in detail (e.g., Kraljic et al. 2012; Scannapieco & Athanassoula 2012; Bonoli et al. 2016). In particular Zana et al. (2018a, b), studying the Eris simulation, focus on the effects of unequal-mass flybys and interactions have found that flybys and mergers have the potential to impact the formation of a bar by inducing a delay in the time of its formation but do not bring about any significant alterations to the overall bar characteristics.
@@
The study of the evolution of barred galaxies and their properties in a statistical framework is now possible also thanks to cosmological hydrodynamic simulations (Vogelsberger et al. 2014a; Schaye et al. 2015; Pillepich et al., 2018b; Nelson et al. 2018). Analysing the EAGLE simulation, Algorry et al. (2017) discover that bars slow down very quickly as they evolve, expanding the inner parts of the dark matter halo. Rosas-Guevara et al. (2020) study massive barred disc galaxies at $z=0$ in the TNG100 simulation (see also Peschken & Lokas 2019; Zhao et al. 2020; Zhou et al. 2020 for Illustris and IllustrisTNG), finding that barred galaxies are less star-forming and more gas poor than unbarred galaxies. Following the evolution of barred galaxies back in time, the authors find that these objects assembled most of their disc components and black holes before bar formation, and earlier than unbarred galaxies. In addition Peschken & Lokas 2019, indicate that flyby interactions are a key mechanism for the formation of bars in high mass disc galaxies in the Illustris simulation whereas the presence of gas in the disc can inhibit the formation of tidally induced bars and weaken existing bars over time.
@@
Regarding the evolution of the bar fraction, Rosas-Guevara et al. (2022) (see also Zana et al. 2022) found that this bar fraction is almost flat as a function of redshift, also finding high redshift bars as early as $z=4$ for the TNG50 simulations and in reasonable agreement with the fraction of the local Universe. The authors also find that if observational biases such as the limited angular resolution, the bar fraction evolution decreases as increasing redshift as seen in observations. Similarly, a decreasing trend was found by Fragkoudi et al. (2021) using the Auriga simulations (Grand et al. 2017). In contrast, Reddish et al. (2022) using the NewHorizon simulation with higher resolution but smaller volume ($10$ Mpc on side), however, found that the bar fraction is smaller than in observations but decreases as increasing redshift. The main reason is that spiral galaxies were dominated by dark matter in the central parts, which do not fulfil the conditions of forming a bar.
@@
Our main goal is to identify the role of the environment in the evolution of the disc galaxies and their bar structures.For that, we follow the evolution of (un)barred disc galaxies ($z=1$) up to $z=0$ and investigate the formation and dissolution of bars and explore if there is connection of the environment in which reside with the formation or disolution of a bar. We make use of the TNG50 simulation (Pillepich et al. 2019; Nelson et al. 2019b) which is the highest resolution simulation run of the TNG project, covering a $51.7$ Mpc region.
@@
The paper is structured as follows. In section 2, we describe the simulation, our parent disc galaxy sample and environment proxies. In section 3, we present the morphological evolution of the disc galaxies up to $z=1$ and their relation with the environment. In section 4, the evolution of the barred galaxies and the fate of their bar structures and in section 5, we investigate the evolution of unbarred galaxies and examine if they form a bar or not. In section 6, we discuss the effects of the environment on the bar fraction. Finally, in section 7, we summarise our findings.
@@ 2.1 TNG Simulations
The subgrid physics of IllustrisTNG hinges on its predecessor, Illustris (Vogelsberger, et al. 2013; Vogelsberger et al. 2014a; Vogelsberger et al. 2014b; Genel et al. 2014; Nelson et al. 2015; Sijacki et al. 2015) with substantial modifications to star formation feedback (winds), the growth of supermassive black holes, Active Galactic Nuclei (AGN) feedback, and stellar evolution and chemical enrichment. A complete description of the improvements in the subgrid physics and calibration process can be found in Pillepich et al., (2018a) and Weinberger et al. (2017). A summary of the improvements concerning Illustris is shown in Table 1 of Pillepich et al., (2018a).
In this work, we focus on the TNG50 simulation (Pillepich et al. 2019; Nelson et al. 2019b), which is the highest resolution simulation that is part of the TNG suite and, at the same time, provides a large enough cosmological volume for studying the statistical properties of galaxies at intermediate masses. The simulation evolves $2160^{3}$ dark matter particles and initial gas cells in a 51.7 comoving Mpc region from $z=127$ down to $z=0$. The mass resolution is $4.5\times 10^{5}\rm M_{\sun}$ for dark matter particles, whereas the mean gas mass resolution is $8.5\times 10^{4}\rm M_{\sun}$. A comparable initial mass is passed down to stellar particles, which subsequently lose mass through stellar evolution. The spatial resolution for collisionless particles (dark matter, stellar, and wind particles) is 575 comoving pc down to $z=1$, after which it remains constant at 288 pc in physical units down to $z=0$. In the case of the gas component, the gravitational softening is adaptive and based on the effective cell radius, down to a minimum value of 72 pc in physical units which is imposed at all times.
Galaxies and their haloes are identified as bound substructures using an FoF and then a SUBFIND algorithm (Springel 2001) and tracked over time by the Sublink merger tree algorithm (Rodriguez-Gomez et al. 2015). Halo masses ($M_{200}$) are defined as all matter within the radius $R_{200}$ for which the inner mean density is $200$ times the critical density. In each FoF halo, the ‘central’ galaxy (subhalo) is the first (most massive) subhalo of each FoF group. The remaining galaxies within the FoF halo are its satellites. The stellar mass of a galaxy is defined as all the stellar matter assigned to the subhalo.
@@ 2.2 Parent disc galaxy sample and identifications of bars
We focused on a subsample of the galaxy evolution, corresponding to $z=1$ disc galaxies contained in the catalogue of Rosas-Guevara et al. (2022) (hereafter RG22). Such galaxies were selected to have a stellar mass larger than $\geq 10^{10}\rm M_{\sun}$ (accounting for more than $10^{5}$ stellar particles), to ensure that the galaxies analysed are well-resolved and being dominated with a disc component ($D/T>0.5$). To identify the disc and bulge components of the galaxies, we use the kinematic decomposition computed in Genel et al. (2015), which is based on Marinacci et al. (2014) and Abadi et al. (2003) and consistent with the selection of the disc sample in Rosas-Guevara et al. (2020) (hereafter RG20) for the TNG100 simulation. Galaxies are first rotated such that the z-axis is located along the direction of the total angular momentum of the stellar component. For each stellar particle within $10\times r_{\rm 50,*}$, where $r_{\rm 50,*}$ is the radius within which 50 per cent of the total stellar mass is contained, the circularity parameter, $\epsilon=J_{z}/J(E)$, is calculated. $J_{z}$ is the specific angular momentum of the particle around the symmetry axis, and $J(E)$ is the maximum specific angular momentum possible at the specific binding energy of each stellar particle. The mass of the stellar disc comprises the stellar particles with $\epsilon\geq 0.7$, while the bulge mass is defined as twice the mass of stellar particles with a circularity parameter $\epsilon<0$. The disc-to-total ($D/T$) ratio is defined as the ratio between the disc stellar mass and the stellar mass enclosed in $10\times r_{50,*}$. We define disc-dominated galaxies as those galaxies with $D/T\geq 0.5$ at $z=1$.
@@
Besides the bulge and disc decomposition described above, we employ the kinematic decomposition mordor computed in Zana et al. (2022) to determine more specific galaxy components. The decomposition is based on the circularity ($\epsilon$) and binding energy ($E$) phase space, where a minimum in $E$ is identified for each galaxy,$E_{\rm cut}$. The following four components can be identified and used throughout the paper:
@@
  • Classical Bulge: is defined as those stellar particles that exhibit the highest degree of binding whose value depends on the $E_{\rm cut}$ of each galaxy and exhibit counterrotation, characterised by a negative value of $\epsilon$ ($<0$). Next, a distribution that is equal to the component of the distribution exhibiting positive circularity is chosen and allocated to the bulge using Monte Carlo sampling.
@@
  • Thin/Cold disc: is defined as those stellar particles that exhibit the highest degree of binding and are not assigned to the bulge but with positive values of $\epsilon$ ($\epsilon>0.7$).
@@
  • Pseudobulge: defined as the remaining stellar particles that exhibit a high degree of binding, but they are not assigned to the bulge nor the thin disc. [^2] Note that this definition of pseudobulge is not necessary to link to the bar and the definition of observed pseudobulge.
@@
  • Thick/Warm disc: is defined as those stellar particles that exhibit a lower degree of binding and are not assigned to the bulge/pseudobulge, but with positive values of $\epsilon$ ($\epsilon>0.7$).

### 07 Feb 2024 v2

@@ Abstract
Context:Stellar bars are non-axisymmetric structures found in over $30$ per cent of massive disc galaxies in the local Universe. The environment could play a significant role in determining whether or not a spiral galaxy is likely to develop a bar.
Aims:We investigate the influence of the environment on the evolution of barred and unbarred disc galaxies with a mass of larger than $10^{10}\rm M_{\sun}$ from $z=1$ down to $z=0$, employing the TNG50 magnetic-hydrodynamical simulation.
Methods:We determined the fraction of barred galaxies that conserve their bar and the fraction of those that lost it by $z=0$. We also estimate the fraction of unbarred galaxies at z=1 that develop a bar at later times. We study the merger histories and the distance of close companions for each category to understand the role of the environment in the evolution of these galaxies.
Results:We find that 49 per cent of $z=1$ disc galaxies undergo a morphological transformation, transitioning into either a lenticular or spheroidal galaxy, while the other $51$ per cent retain the large disc shape. The morphological alteration is mostly influenced by the environment. Lenticular and spheroidal galaxies tend to exist in denser environments and have more frequent mergers compared to disc galaxies. We find that bars are stable after they have formed, as over half of the barred galaxies (60.2 per cent) retain the bar structure and have experienced fewer mergers compared to those galaxies that lose their bars (5.6 per cent). These latter galaxies start with weaker and shorter bars at $z=1,$ are influenced by tidal interactions, and are frequently observed in more populated areas. Additionally, our study reveals that less than 20 per cent of unbarred galaxies will never develop a bar and exhibit the quietest merger history. Unbarred galaxies that undergo bar formation after $z=1$ more frequently experience a merger event. Furthermore, tidal interactions with a close companion may account for bar formation in at least one-third of these instances.
@@ 1 Introduction
Stellar bars are non-axisymmetric structures that are found in the central regions of the majority of disc galaxies. They are present in over $30$ per cent of massive disc galaxies with a stellar mass of $10^{10}M_{\odot}$ in the local Universe (e.g. Sellwood & Wilkinson 1993; Masters et al. 2011; Gavazzi et al. 2015). These bars are believed to be drivers of the secular evolution of galaxies, as they are efficient in redistributing gas, stars, and dark matter towards the central regions of the galaxy (e.g. Debattista et al. 2004; Athanassoula et al. 2005).
The presence or absence of bars in galaxies can reveal crucial information about the assembly history of these galaxies. One effective way to investigate this is by studying the evolution of bar fractions as a function of various galaxy properties, such as stellar mass, gas content, or disc structure. Local Universe studies have shown that bars are more frequently found in massive, gas-poor, and red galaxies than in their blue, gas-rich, and star-forming counterparts (e.g. Barazza et al. 2008; Masters et al. 2012; Gavazzi et al. 2015; Consolandi 2016; Cervantes-Sodi 2017). Moreover, the redshift dependence of the bar fraction can provide additional insights into the evolution of bars and their host galaxies.
@@
Explorations of whether or not the bar fraction is increasing or decreasing can be used to study the lifetimes of bar structures providing useful information about whether these structures are stable in time or disappear through secular processes (e.g. Aguerri, Méndez-Abreu, & Corsini 2009) or through external processes (e.g. Zana et al. 2018b; Peschken & Lokas 2019; Méndez-Abreu, Costantin, & Kruk 2023). This topic is hotly debated. The secular process of forming and dissolving a stellar bar can be complex and takes place over long timescales. Bars can be formed spontaneously via secular evolution as a result of global, non-axisymmetric instabilities (e.g. Athanassoula 2003; Mendez-Abreu et al. 2012; Gavazzi et al. 2015). On the other hand, the combined effects of central mass concentrations (CMCs) and gravity torques could make galactic bars transient features, with lifetimes of 1-2 Gyr in typical Sb-Sc galaxies (Bournaud, Combes, & Semelin 2005).
@@
The formation of new stars can lead to feedback processes that release energy and momentum that can disrupt the bar structure and gradually weaken and dissolve it (e.g. Zana et al. 2019). Vertical resonances can also arise within the bar, causing stars to oscillate vertically within the bar. This can lead to the formation of a central mass concentration, which can destabilise the bar and cause it to dissolve (e.g. Martinez-Valpuesta, Shlosman, & Heller 2006).
@@
In the case of external processes, galactic mergers and interactions can play a significant role in the formation and disruption of stellar bars in galaxies (e.g. Elmegreen, Elmegreen, & Bellin 1990; Berentzen et al. 2004; Łokas et al. 2016; Zana et al. 2018a, b; Łokas 2019). The specific conditions required for a galactic merger or interaction to form or dissolve a bar include the mass ratio of the merging galaxies, and the phase of the bar and of the orbit, because the strength of the perturbation can be a deciding factor in whether or not the central part of the galaxy can form a bar (Elmegreen, Elmegreen, & Bellin 1990). The orbit of the merging galaxies can be another condition; a direct collision between two galaxies can cause significant perturbations that can lead to the formation or dissolution of a bar (Berentzen et al. 2004). Moreover, prograde encounters (when the angular momenta of the two galaxies are aligned) have a much more dramatic effect on the galactic structure than retrograde ones (when the angular momenta point in opposite directions). Prograde encounters lead to the formation of long and narrow tidal arms and tidally induced bars in galaxies orbiting a larger host, but retrograde ones do not (Łokas 2018). The gas content of the merging galaxies also contributes to the dissolution of a bar. If the merging galaxies have a significant amount of gas for instance, tidal interactions due to minor mergers can weaken the central stellar bar in the merger remnant (Ghosh et al. 2021).
@@
Tidal forces due to dense environments or close companions also have a significant impact on the formation or dissolution of bars. Łokas et al. (2014) and Łokas et al. (2016) study the formation of a bar in galaxy encounters with a Milky-way galaxy and dwarf galaxy embedded in dense environments using N-body simulations. The authors find that the characteristics of the bars undergo temporal variations and are subject to the influence of the magnitude of the tidal force encountered. The formation of bars in galaxies is observed to occur at earlier stages and the bar itself is of greater strength and length in galaxies that experience a higher degree of tidal force exerted by the surrounding cluster.
@@
It has very recently become possible to study the external and internal processes of galaxies in detail with zoomed-in simulations (e.g. Kraljic et al. 2012; Scannapieco & Athanassoula 2012; Bonoli et al. 2016). In particular, Zana et al. (2018a, b), studying the Eris simulation, focus on the effects of unequal-mass flybys and interactions. These authors find that flybys and mergers have the potential to impact the formation of a bar by inducing a delay in the time of its formation but do not bring about any significant alterations to the overall bar characteristics.
@@
The study of the evolution of barred galaxies and their properties in a statistical framework is now also possible thanks to cosmological hydrodynamic simulations (Vogelsberger et al. 2014a; Schaye et al. 2015; Pillepich et al., 2018b; Nelson et al. 2018). Analysing the EAGLE simulation, Algorry et al. (2017) discovered that bars slow down very quickly as they evolve, expanding the inner parts of the dark matter halo. Rosas-Guevara et al. (2020) studied massive barred disc galaxies at $z=0$ in the TNG100 simulation (see also Peschken & Lokas 2019; Zhao et al. 2020; Zhou et al. 2020 for Illustris and IllustrisTNG), finding that barred galaxies are less star forming and more gas poor than unbarred galaxies. Following the evolution of barred galaxies back in time, the authors find that these objects assembled most of their disc components and black holes before bar formation, and did this earlier than unbarred galaxies. In addition, Peschken & Lokas 2019, indicate that flyby interactions are a key mechanism in the formation of bars in high-mass disc galaxies in the Illustris simulation, whereas the presence of gas in the disc can inhibit the formation of tidally induced bars and weaken existing bars over time.
@@
Regarding the evolution of the bar fraction, Rosas-Guevara et al. (2022) (see also Zana et al. 2022) found that this bar fraction is almost flat as a function of redshift, also finding high-redshift bars as early as $z=4$ for the TNG50 simulations and in reasonable agreement with the fraction of the local Universe. The authors also find that if there are observational biases, such as those caused by limited angular resolution, the bar fraction evolution decreases with increasing redshift, as seen in observations. Similarly, a decreasing trend was found by Fragkoudi et al. (2021) using the Auriga simulations (Grand et al. 2017). In contrast, Reddish et al. (2022), using the NewHorizon simulation with higher resolution but smaller volume ($10$ Mpc on side), found that the bar fraction is smaller than in observations but decreases with increasing redshift. The main reason for this was found to be that spiral galaxies are dominated by dark matter in the central parts, which do not fulfil the conditions of forming a bar.
@@
The main goal of the present study is to identify the role of the environment in the evolution of disc galaxies and their bar structures. To this end, we follow the evolution of barred and unbarred disc galaxies ($z=1$) up to $z=0$ and investigate the formation and dissolution of their bars. We also explore whether or not there is a connection between the environment in which a galaxy resides and the likelihood that a bar will form or dissolve therein. We make use of the TNG50 simulation (Pillepich et al. 2019; Nelson et al. 2019b), which is the highest-resolution simulation run of the TNG project, covering a $51.7$ Mpc region.
@@
The paper is structured as follows. In section 2, we describe the simulation, our parent disc galaxy sample, and environment proxies. In section 3, we present the morphological evolution of the disc galaxies up to $z=1$ and their relation with the environment. In section 4, we study the evolution of the barred galaxies and the fate of their bar structures. In section 5, we investigate the evolution of unbarred galaxies and examine whether or not they form a bar. In section 6, we discuss the effects of environment on bar fraction. Finally, in section 7, we summarise our findings.
@@ 2.1 TNG simulations
The subgrid physics of IllustrisTNG is based on its predecessor, Illustris (Vogelsberger, et al. 2013; Vogelsberger et al. 2014a; Vogelsberger et al. 2014b; Genel et al. 2014; Nelson et al. 2015; Sijacki et al. 2015), with substantial modifications to star formation feedback (winds), the growth of supermassive black holes, active galactic nucleus (AGN) feedback, and stellar evolution and chemical enrichment. A complete description of the improvements made to the subgrid physics and calibration process can be found in Pillepich et al., (2018a) and Weinberger et al. (2017). A summary of the improvements concerning Illustris is shown in Table 1 of Pillepich et al., (2018a).
In this work, we focus on the TNG50 simulation (Pillepich et al. 2019; Nelson et al. 2019b), which is the highest-resolution simulation that is part of the TNG suite and, at the same time, provides a sufficiently large cosmological volume to study the statistical properties of galaxies at intermediate masses. The simulation evolves $2160^{3}$ dark matter particles and initial gas cells in a 51.7 comoving 1 Mpc region from $z=127$ down to $z=0$. The mass resolution is $4.5\times 10^{5}\rm M_{\sun}$ for dark matter particles, whereas the mean gas mass resolution is $8.5\times 10^{4}\rm M_{\sun}$. A comparable initial mass is passed down to stellar particles, which subsequently lose mass through stellar evolution. The spatial resolution for collisionless particles (dark matter, stellar, and wind particles) is 575 comoving pc down to $z=1$, after which it remains constant at 288 pc in physical units down to $z=0$. In the case of the gas component, the gravitational softening is adaptive and based on the effective cell radius down to a minimum value of 72 pc in physical units, which is imposed at all times.
Galaxies and their haloes are identified as bound substructures using a friends-of-friends (FoF) and then a SUBFIND algorithm (Springel 2001) and are then tracked over time by the Sublink merger tree algorithm (Rodriguez-Gomez et al. 2015). Halo masses ($M_{200}$) are defined as all matter within the radius $R_{200}$ for which the inner mean density is $200$ times the critical density. In each FoF halo, the ‘central’ galaxy (subhalo) is the first (most massive) subhalo of each FoF group. The remaining galaxies within the FoF halo are its satellites. The stellar mass of a galaxy is defined as all the stellar matter assigned to the subhalo.
@@ 2.2 Parent disc galaxy sample and identifications of bars
We focused on a subsample of the galaxy evolution, corresponding to $z=1$ disc galaxies contained in the catalogue of Rosas-Guevara et al. 2022, hereafter RG22. Such galaxies were selected to have a stellar mass of $\geq 10^{10}\rm M_{\sun}$ (accounting for more than $10^{5}$ stellar particles) to ensure that the galaxies analysed are well resolved and are dominated by a disc component ($D/T>0.5$). To identify the disc and bulge components of the galaxies, we use the kinematic decomposition computed in Genel et al. (2015), which is based on Marinacci et al. (2014) and Abadi et al. (2003) and is consistent with the selection of the disc sample in Rosas-Guevara et al. 2020, hereafter RG20 for the TNG100 simulation. Galaxies are first rotated such that the z-axis is located along the direction of the total angular momentum of the stellar component. For each stellar particle within $10\times r_{\rm 50,*}$, where $r_{\rm 50,*}$ is the radius within which 50 per cent of the total stellar mass is contained, the circularity parameter, $\epsilon=J_{z}/J(E)$, is calculated. $J_{z}$ is the specific angular momentum of the particle around the symmetry axis, and $J(E)$ is the maximum specific angular momentum possible at the specific binding energy of each stellar particle. The mass of the stellar disc comprises the stellar particles with $\epsilon\geq 0.7$, while the bulge mass is defined as twice the mass of the stellar particles, with a circularity parameter $\epsilon<0$. The disc-to-total ($D/T$) mass ratio is defined as the ratio between the disc stellar mass and the stellar mass enclosed in $10\times r_{50,*}$. We define disc-dominated galaxies as those galaxies with $D/T\geq 0.5$ at $z=1$.
@@
In addition to the bulge and disc decomposition described above, we employ the kinematic decomposition mordor computed in Zana et al. (2022) [^2] https://github.com/thanatom/mordor to determine more specific galaxy components. The decomposition is based on the circularity ($\epsilon$) and binding energy ($E$) phase space, where a minimum in $E$ is identified for each galaxy, $E_{\rm cut}$. The following four components can be identified and are used throughout the paper:
@@
  • Classical bulge: defined as those stellar particles that exhibit the highest degree of binding, the value of which depends on the $E_{\rm cut}$ of each galaxy, and exhibit counter-rotation characterised by a negative value of $\epsilon$ ($<0$). Next, a distribution that is equal to the component of the distribution exhibiting positive circularity is chosen and allocated to the bulge using Monte Carlo sampling.
@@
  • Thin/cold disc: defined as those stellar particles that exhibit the highest degree of binding and are not assigned to the bulge but have positive values of $\epsilon$ ($\epsilon>0.7$).
@@
  • Pseudobulge: defined as the remaining stellar particles that exhibit a high degree of binding, but are not assigned to the bulge or the thin disc. [^3] We note that this definition of pseudobulge is not necessary to link to the bar and the definition of observed pseudobulge.
@@
  • Thick/warm disc: defined as those stellar particles that exhibit a lower degree of binding and are not assigned to the bulge or pseudobulge, but with positive values of $\epsilon$ ($\epsilon>0.7$).